# DRCT: Saving Image Super-Resolution away from Information Bottleneck

### v1

@@ 1 Introduction
Figure 1: The feature map intensity on various benchmark datasets. We observed that feature map intensities decrease sharply at deeper network levels, indicating potential information loss. We propose DRCT, stabilizes this by enhancing receptive fields and adding dense-connections within residual blocks to mitigate information bottlenecks, thereby improving performance with a simpler model.
@@
Inspired by a series of works by Wang et al., such as the YOLO family [42, 39], CSPNet [40], and ELAN [41], we consider that network architectures based on SwinIR, despite significantly enlarging the receptive field through shift-window attention to address the issue of insufficient receptive fields in CNNs, are prone to gradient bottlenecks due to the loss of spatial information as network depth increases. This implicitly constrains the model’s performance.
@@ 3.2 Spatial Information Vanish in Super-resolution
$GI=\sum^{L}_{l=1}(\left\|\Delta G_{min}^{l}\right\|+\left\|\Delta G_{max}^{l}\right\|)\propto\text{Information loss}$ (4)
@@ 4.3 Same-task Progressive Training
and finally use L2 loss to eliminate singular pixels and artifacts, therefore archiving state-of-the-arts performance.
@@ 5.2 Implementation Details
The training process can be structured into three distinct phases, as ullustrated as Section 4-3. (1) pre-trained on ImageNet, (2) optimize the model on the given dataset, (3) L2-loss for PSNR enhancement. Throughout the training process, we use the Adam optimizer with $\beta_{1}$ = $0.9$, and $\beta_{2}$ = $0.999$ and train for $800k$ iterations in each stage. The learning rate is set to $42e-$ 4, the multi-step learning scheduler is also used. The learning rate is halved at the $300k$, $500k$, $650k$, $700k$, $750k$ iterations respectively. Weight decay is not applied and the batch-size is set to $32$. In the architecture of DRCT, the configuration of depth and width is maintained identically to that of SwinIR. To elaborate, both the number of RDFEG and SDRCB units are established at 6, and the channel number of intermediate feature maps is designated as 180. The attention head number and window size are set to 6 and 16 for window-based multi head self-attention (W-MSA). In terms of data preparation, HR patches with dimensions of $256$$\times$ $256$ pixels were extracted from the HR images. To improve the generalizability, we apply random horizontal flips and rotation augmentation.
@@ 5.3 Quantitative Results
For the evaluation, we use full RGB channels and ignore the $(2\times\text{scale})$ pixels from the border. PSNR and SSIM metrics are used to evaluation criteria. Table 1 presents the quantitative comparison of our approach and the state-of-the-art methods, including EDSR [25], RCAN [52], SAN [8], IGN [56], HAN [32], NLSN [31], SwinIR [24], as well as approaches using ImageNet pre-training, such as IPT [3], EDT [22] and HAT [5]. We can see that our method outperforms the other methods significantly on all benchmark datasets. In addition, the DRCT-L can bring further improvement and greatly expands the performance upper bound of this task. DRCT with fewer parameters and similar computation can also significantly outperforms the state-of-the-art method HAT.
@@ 5.4 Visual Comparison
The visual comparisons displayed in Figure 5. For the selected images from the Urban100 dataset [17], DRCT is effective in restoring structures, whereas other methods suffer from notably blurry effects. The visual results demonstrate the superiority of our approach.
@@ 5.5 Model Complexity
To demonstrate the potential of our proposed method, we conducted further analysis on model complexity, with the results illustrated in Figure 7. From the figure, we can observe that the performance curves of the HAT and SwinIR models are approaching horizontal lines, suggesting that their model performance is nearing a bottleneck. Enhancing performance may require significantly more complex models or a greater number of parameters, potentially leading to prohibitive costs and practical implementation challenges. In contrast, our DRCT-L model outperforms HAT-L by 0.14dB in PSNR while saving 33% of model parameters. This indicates that stabilizing the training and inference processes within residual blocks through a dense-group is a viable strategy to prevent SISR models from encountering information bottlenecks.

### v2

@@ 1 Introduction
Figure 1: The feature map intensity on various benchmark datasets. We observed that feature map intensities decrease sharply at deeper network levels, indicating potential information loss. We propose DRCT, stabilizes this by enhancing receptive fields and adding dense-connections within residual blocks to mitigate information bottlenecks, thereby improving performance with a simpler model design.
@@
Inspired by a series of works by Wang et al., such as the YOLO-family [42, 39], CSPNet [40], and ELAN [41], we consider that network architectures based on SwinIR, despite significantly enlarging the receptive field through shift-window attention to address the issue of insufficient receptive fields in CNNs, are prone to gradient bottlenecks due to the loss of spatial information as network depth increases. This implicitly constrains the model’s performance.
@@ 3.2 Spatial Information Vanish in Super-resolution
$GI=\sum^{L}_{l=1}(\left\|\Delta G_{min}^{l}\right\|+\left\|\Delta G_{max}^{l}\right\|)$ (4)
@@ 4.3 Same-task Progressive Training
and finally use L2 loss to eliminate singular pixels and artifacts, therefore further archiving greater performance on PSNR metrics.
@@ 5.2 Implementation Details
The training process can be structured into three distinct phases, as ullustrated as Section 4-3. (1) pre-trained on ImageNet [9], (2) optimize the model on the given dataset, (3) L2-loss for PSNR enhancement. Throughout the training process, we use the Adam optimizer with $\beta_{1}$ = $0.9$, and $\beta_{2}$ = $0.999$ and train for $800k$ iterations in each stage. The learning rate is set to $42e-$ 4, the multi-step learning scheduler is also used. The learning rate is halved at the $300k$, $500k$, $650k$, $700k$, $750k$ iterations respectively. Weight decay is not applied and the batch-size is set to $32$. In the architecture of DRCT, the configuration of depth and width is maintained identically to that of SwinIR. To elaborate, both the number of RDFEG and SDRCB units are established at 6, and the channel number of intermediate feature maps is designated as 180. The attention head number and window size are set to 6 and 16 for window-based multi head self-attention (W-MSA). In terms of data preparation, HR patches with dimensions of $256$$\times$ $256$ pixels were extracted from the HR images. To improve the generalizability, we apply random horizontal flips and rotation augmentation.
@@ 5.3 Quantitative Results
For the evaluation, we use full RGB channels and ignore the $(2\times\text{scale})$ pixels from the border. PSNR and SSIM metrics are used to evaluation criteria. Table 1 presents the quantitative comparison of our approach and the state-of-the-art methods, including EDSR [25], RCAN [52], SAN [8], IGN [56], HAN [32], NLSN [31], SwinIR [24], CAT-A [6], DAT [7], as well as approaches using ImageNet pre-training, such as IPT [3], EDT [22] and HAT [5]. We can see that our method outperforms the other methods significantly on all benchmark datasets. In addition, the DRCT-L can bring further improvement and greatly expands the performance upper-bound of this task. Even with fewer model parameters and computational requirements, DRCT is also significantly greater than the state-of-the-art methods.
@@ 5.4 Visual Comparison
The visual comparisons displayed in Figure 5. For the selected images from the Urban100 [17], DRCT is effective in restoring structures, whereas other methods suffer from notably blurry effects. The visual results demonstrate the superiority of our approach.
@@
There remains significant room for improvement in the way information loss is measured for SISR models. We conducted a detailed analysis of the G-index using ’image_092’ from Urban100, as shown in Figure 6. DRCT managed to achieve a lower score on the GI, indicating that the changes in the network’s feature maps are relatively stable. In contrast, The GI for HAT was significantly higher than the other two, which suggests that a lower GI value can effectively denote lesser information loss, whereas a higher GI might imply more drastic changes in feature maps. However, this doesn’t necessarily correlate directly with the effectiveness of measuring information loss.
@@ 5.5 Model Complexity
To demonstrate the potential of our proposed method, we conducted further analysis on model complexity, with the results illustrated in Figure 7. From the figure, we can observe that the performance curves of the HAT and SwinIR models are approaching horizontal lines, suggesting that their model performance is nearing a bottleneck. Enhancing performance may require significantly more complex models or a greater number of parameters, potentially leading to prohibitive costs and practical implementation challenges. In contrast, our DRCT-L model outperforms HAT-L while saving 33% of model parameters. This indicates that stabilizing the training and inference processes within residual blocks through a dense-group is a viable strategy to prevent SISR models from encountering information bottlenecks.